a man walks 3.00 mile due east, then turns and walks 2.00 mi due north. how far is he from the starting point
Answer #1

Displacement,

D^2 = Dx^2 + Dy^2

D^2 = 3^2 + 2^2

D = 3.606 miles

========

Comment in case any doubt, will reply for sure.. Goodluck